10. A binomial distribution describes a random variable, X, that has 4 possible outcomes. Outcome 1 has a probability of 25%, Outcome 2 has a probability of 15%, Outcome 3 has a probability of 20%. What is the probability of Outcome 4?    (5 pts)
Answer #1

Solution:

Given in the question

P(Outcome 1) = 0.25

P(Outcome 2) = 0.15

P(Outcome 3) = 0.2

So we need to find out P(Outcome 4) = ?

From binomial distribution theorem we can say that total probability is 1

So P(Outcome 4) = 1-P(Outcome 1)-P(Outcome2) - P(Outcome3) = 1-0.25-0.15-0.2 = 1-0.6 = 0.4

So P(Outcome4) = 0.4

So binomial distribution table can be constructed as

X P(X)

1 0.25

2 0.15

3 0.2

4 0.4